I need to know how to solve this by substitution.
x=y-6
2x+3y=33

Answer by nerdybill(7384)

x=y-6 (equation 1)
2x+3y=33 (equation 2)
.
substitute definition of x from equation 1 into equation 2 and solve for y:
Start with equation 2:
2x+3y=33
do our substitution:
2(y-6)+3y=33
2y-12+3y=33
-12+5y=33
5y=45
y = 9
.
substitute above into equation 1:
x=y-6
x=9-6
x = 3
.
Solution:
x = 3
and
y = 9